artlover03:

Kyle and Myles wanted to buy the soccer team food after the game. Kyle bought 5 hotdogs and 3 French fries, which costs him $14.50. Myles bought 4 hotdogs and 2 French fries, which totaled as $11.00. How much does 1 hotdog cost? Walkthrough your answer by showing your work

snowflake0531:

do you know how to make the equations for this?

artlover03:

no i dont :(

artlover03:

i will probably catch on once i get some help

snowflake0531:

okay, so, let's first name the variables which two things do you think will be our x adn y

artlover03:

x= hotdogs y=fries ??

snowflake0531:

yes lol so, here it says 5 hot dogs and 3 french fries costs 14.50 dollars how would you put that into an equation

artlover03:

5x+3y= 14.50 omg i see it now

snowflake0531:

loll, ye and for teh second one?

artlover03:

4x+2y=11.00

snowflake0531:

yesssss so now that you got the equations 5x+3y=14.5 and 4x+2y=11 do you want to use the substitution method, elimination, or graphing (graphing is easier, but teachers want you to be able to solve it yourself, so if you know how to do it, just graph it lol)

artlover03:

actually im lost, how do i isolate the variable

snowflake0531:

ohhh, okay then we'll do the substitution method do you want to first solve for x or y?

artlover03:

x

snowflake0531:

okay, then for the first equation, we first isolate the y 5x+3y=14.5 3y= -5x + 14.5 y= -5/3 x + 29/6 this is going to be messy -_- but do you follow?

snowflake0531:

so in regular stuff like x+ y = __ and x - y= ___ first isolate the y, to y=____ then put the y value in, and substitute it into the second equation, that's why it's called the substitution method

snowflake0531:

and because it's going to be more confusing to do substitution in this than elimination, you should learn the elimination in this one lol 5x+3y=14.5 and 4x+2y=11 since you're solving for x, you want to eliminate the y so, multiply the top equation, all, by 2, and the bottom equation, all, by 3

artlover03:

okay im up to y= -5/3 -*y* how did you get 29/6

snowflake0531:

bc i did 14.5/3 and tehn did 145/30 and then simplified lol and bc of the weird numbers... that's why substitution for this one will be too confusing-_- lol

artlover03:

i got -4.83 repeating is that right? for 14,50/3

snowflake0531:

....no 14.5/3... how'd you get a negative lol

snowflake0531:

also, bc it keeps on going on, you keep it in a fraction

artlover03:

because you moved the 3y to one side by subtracting.. then to get it by itself you divude by -3

snowflake0531:

but the 14.5 is still positive

artlover03:

how? dont you divide the entire equatioin by -3? so it would be 5x+3y=14.50 -3y= 5x+14.50 then divide all of it by -3 to get the y variable by itself y= -5/3 -14.50/3 or -4.83

snowflake0531:

O-O i am confused at what you're doing 5x+3y=14.5 3y = -5x + 14.5 ^^ stays positive

artlover03:

oh actually i see how to do it: i subsitute 0 for a variable: 5x+3(0)=14.50 5x=14.50 x=2.90 for Kyles purchase 4x+2(0)=11.00 4x=11.00 x= 2.75 for myles

snowflake0531:

i- no lol

artlover03:

no im wrong this is the problem (i changed the names)

1 attachment
snowflake0531:

ummm, how about we just graph it graph the two equations that you wrote in desmos and fidn the coordinate in which they intersect (x,y) x is hot dogs y is french fries

artlover03:

i didnt write in desmos?

snowflake0531:

ik you didn't

artlover03:

... im lost

snowflake0531:

the point in intersection (x,y) in this case (2,1.5) so, x is 2, which is hot dogs, so 2 dollars y is 1.5 dolalrs, french fries

artlover03:

wait how you know its 2 dollars for the hotdogs?? im trying to solve for that? @jhonyy9

snowflake0531:

because i graphed it... and when you put in two linear equations... teh point in intersection is where you find the x and teh y

artlover03:

for kyle, the total for both the 5 hotdogs and 3 fries was 14.50. For Myles, the total for 4 hotdogs and 2 fries was 11.00. my question wants to know the price for one. im dont understand how the graph answer matches my answer choices'

jhonyy9:

5x +3y = 14,50 4x +2y = 11 solve this system for x and y

jhonyy9:

@artlover03 is offline

snowflake0531:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @artlover03 for kyle, the total for both the 5 hotdogs and 3 fries was 14.50. For Myles, the total for 4 hotdogs and 2 fries was 11.00. my question wants to know the price for one. im dont understand how the graph answer matches my answer choices' \(\color{#0cbb34}{\text{End of Quote}}\) bc hot dog is x.... french fries is y... x is 2, y is 1.5 so, 2 dollars and 1.50 dollars

jhonyy9:

ok but then why he/she asked it - posted it above newly ?

snowflake0531:

?

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @artlover03 for kyle, the total for both the 5 hotdogs and 3 fries was 14.50. For Myles, the total for 4 hotdogs and 2 fries was 11.00. my question wants to know the price for one. im dont understand how the graph answer matches my answer choices' \(\color{#0cbb34}{\text{End of Quote}}\) here is